### v1

Abstract
Accurate and efficient seismic response prediction is essential for the design of resilient structures. While the Finite Element Method (FEM) remains the standard for nonlinear seismic analysis, its high computational demands limit its scalability and real-time applicability. Recent developments in deep learning—particularly Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), and Long Short-Term Memory (LSTM) models—have shown promise in reducing the computational cost of the nonlinear seismic analysis of structures. However, these data-driven models often struggle to generalize and capture the underlying physics, leading to reduced reliability. We propose a novel Physics-Informed U-Net-LSTM framework that integrates physical laws with deep learning to enhance both accuracy and efficiency. By embedding domain-specific constraints into the learning process, the proposed model achieves improved predictive performance over conventional Machine Learning (ML) architectures. This hybrid approach bridges the gap between purely data-driven methods and physics-based modeling, offering a robust and computationally efficient alternative for seismic response prediction of structures.
keywords:
[…]
{graphicalabstract}[Uncaptioned image]{highlights}
A novel PhyULSTM framework is developed for data-driven structural seismic modeling.
Physics laws in loss guide learning, prevent overfitting, and improve generalization.
The model can predict seismic response without physical constraints as well.
1 Introduction
The increasing number and intensity of natural disasters such as earthquakes, typhoons, and tsunamis have heightened the global demand for resilient and adaptive infrastructure systems. Among these hazards, seismic events pose particularly severe threats to structural integrity, leading to widespread damage and loss of life. Accurate prediction of structural response under seismic loading is thus a cornerstone of resilient design. Conventional numerical methods, including the Newmark-$\beta$ [newmark1959method], KR-$\alpha$ [kolay2014development], and the finite element method (FEM)[reddy1993introduction, huebner2001finite, zienkiewicz2005finite], have long been the foundation for modeling dynamic behavior of structures. These techniques are reliable and grounded in physics, yet they are computationally intensive, especially when applied to nonlinear analysis of a full-scale structure. Their high computational cost limits real-time applications and scalability.
In recent years, machine learning (ML) and artificial intelligence (AI) have emerged as promising alternatives or supplements to traditional physics-based modeling. A range of ML techniques,including support vector machines (SVM) [yinfeng2008nonlinear, segura2020metamodel, gharehbaghi2020estimating], multilayer perceptrons (MLP) [NN_ref1, NN_ref2, NN_ref3, NN_ref4, NN_ref5, NN_ref6, NN_ref7, NN_ref8, NN_ref9, lagaros2012neural], deep neural networks [vaidyanathan2005artificial, kim2019response, kim2020probabilistic], long short-term memory networks (LSTMs) [zhang2019deep], and convolutional neural networks (CNNs) [wu2019deep], have been employed to develop surrogate models for rapid prediction of structural response. These approaches reduce computational cost while offering acceptable levels of accuracy. RNNs and LSTMs, in particular, are often adopted for modeling temporal dynamics in structural systems due to their ability to learn long-term dependencies in time series data [graves2012long]. Meanwhile, CNNs with dilated filters have shown promise for time-series modeling by enabling access to long-range dependencies without recurrence [bai2018empirical, oord2016wavenet]. These models have also been successfully applied in domains such as ECG classification, structural health monitoring, continuum damage mechanics and anomaly detection [bird2020cross, tang2019convolutional, danilczyk2021smart, zhao2019speech, pantidis2023integrated, pantidis2024fenn].
Building upon these advances, physics-informed machine learning has gained traction within earthquake engineering and structural dynamics community, where it addresses critical challenges such as data scarcity and limited model generalization. By embedding governing physical laws directly into ML architectures, physics-informed frameworks enhance predictive accuracy and ensure physical consistency. For instance, physics-informed LSTM frameworks (e.g., PhyLSTM2, PhyLSTM3)[phylstm] estimate key variables like displacement and restoring force by incorporating these laws into the loss function during training. Similarly, PhyCNN [phycnn] integrates domain knowledge with deep CNNs to predict nonlinear seismic responses through a data-driven approach. These innovative methods have been applied across diverse scenarios, including subway station dynamics [huang2021deep], vehicle–bridge interactions [li2021dynamic], and dam behavior prediction using CNN–GRU architectures [li2022new]. More recently, PI-LSTM has been introduced and benchmarked against PhyCNN and PhyLSTM, further demonstrating the potential of physics-guided ML architectures for structural response prediction [liu2023pi]. Recent works [raissi2018deep, raissi2019physics, sun2020surrogate, zhu2019physics] have highlighted the importance of incorporating physical constraints—such as ordinary and partial differential equations and boundary conditions—directly into deep learning models to improve stability and accuracy even when data is limited.
Despite the progress made, key limitations persist. LSTMs and RNNs, while effective in modeling temporal dependencies, suffer from long training times and are prone to overfitting, particularly in low-data regimes [bengio1994learning, trinh2018learning]. Deep CNNs, although powerful, require substantial computational resources when applied to long time series, and their performance degrades under large plastic deformations [kuo2024gnn]. Moreover, capturing long-range interactions and nonlinearities in dynamic systems remains a fundamental challenge. To overcome these challenges, we propose a novel Physics-Informed U-net–LSTM architecture for the data-driven modeling of structural seismic response. The model incorporates a causal one-dimensional U-Net, developed leveraging the “slow feature” hypothesis [wiskott2002slow], to extract hierarchical temporal features across multiple scales while preserving temporal causality. These features are then passed to an LSTM network, enabling efficient sequence modeling and dynamic prediction. Importantly, physical knowledge—expressed through governing equations of motion—is embedded into the loss function, guiding the model to honor the laws of mechanics even in data-scarce scenarios. The framework is validated on both numerical and experimental benchmarks and is compared with PhyCNN framework. Even without knowledge of the structural system, it can still predict the response only by utilizing the data loss in a purely data-driven manner. Results demonstrate that the proposed approach offers superior accuracy, robustness, and generalization, especially under conditions of strong nonlinearity, making it a viable surrogate for seismic response prediction of complex structural systems.
2 Physics-Informed U-net-LSTM Network
In this section, we present the PhyULSTM framework for surrogate modeling of nonlinear dynamic structural systems subjected to seismic ground motion. PhyULSTM consists of three core components: a 1D U-Net, a deep LSTM network, and a graph-based tensor differentiator. The proposed 1D U-Net, a tailored variant of the standard U-Net architecture, is designed to work synergistically with deep LSTM networks for time-series analysis. It has potential to alleviate common limitations of traditional RNNs and LSTMs—such as difficulty in capturing long-range temporal dependencies and susceptibility to overfitting—especially when processing high-resolution seismic inputs like ground acceleration. This is achieved by leveraging the ‘Slow Feature’ hypothesis [wiskott2002slow], which posits that the most informative features in temporal sequences evolve slowly over time, even when raw signals exhibit rapid fluctuations. By extracting such temporally stable features across multiple time scales and applying causal convolutions—which restrict each prediction to depend only on current and past inputs—the U-Net enhances the LSTM’s ability to learn long-term dependencies within an auto-regressive framework. Additionally, physical knowledge is explicitly incorporated into the model through a graph-based tensor differentiator (based on the finite difference method), as introduced by Zhang et al. [phycnn] in their PhyCNN framework. This component computes derivatives of the state-space outputs to construct a physics-informed loss term derived from the system’s governing equations. The hybrid U-Net–LSTM model is trained on seismic input–output datasets—obtained either through numerical simulation or sensor measurements—and subsequently deployed as a surrogate model for predicting structural responses under future seismic excitations. A detailed illustration of the PhyULSTM framework is presented in Figure 1.
To demonstrate the PhyULSTM framework, we consider a general dynamic structural system subjected to ground motion excitation, governed by the equation:
$M\ddot{x}(t)+h(t)=-M\Gamma\ddot{x}_{g}(t),$ (1)
where $M$ is the mass matrix; $x(t)$ and $\dot{x}(t)$ denote the relative displacement and velocity with respect to the ground; $\ddot{x}_{g}(t)$ is the ground acceleration vector; $\Gamma$ represents the force distribution vector; and $h(t)$ denotes the latent generalized restoring force vector, which is modeled as a function of the ground acceleration $\ddot{x}_{g}(t)$ and the structural acceleration with respect to the ground $\ddot{x}(t)$. Normalizing Eq. (1) by $M$, the governing equation can be rewritten as
$f:\ddot{x}(t)+g(t)+\Gamma\ddot{x}_{g}(t)\rightarrow 0,$ (2)
where $g(t)=M^{-1}h(t)$ is the mass-normalized restoring force. The explicit form of $g(t)$ is assumed unknown and is therefore learned implicitly by the proposed PhyULSTM framework.
Figure 1: The proposed Physics-Informed U-net-LSTM Network (PhyULSTM) for time-series modeling. The 1D U-net first receives the ground acceleration data as input and extracts features at multiple time scales and adapts it to our auto-regressive scenario by ensuring that the convolutions are causal. The outputs from the U-net are then fed to the deep LSTM network that maps the temporal feature maps to the corresponding output space. The outputs are state space variables z(t){z}(t) = $\{x(t),\dot{x}(t),g(t)\}$. Available physics knowledge is incorporated directly into the loss function. In addition to it, a graph-based tensor differentiator using the central finite difference method developed by Zhang et.al in their PhyCNN framework is implemented to calculate the derivative of state space outputs ${z_{t}}(t)$ = $\{x_{t}(t),\dot{x}_{t}(t),g_{t}(t)\}$ to construct the physics loss from the governing equation, where the subscript t represents the derivative of the state with respect to time. By optimizing the network hyperparameters $\theta=\{W_{\theta},b_{\theta}\}$, PhyULSTM can interpret the measurement data (e.g. $\{x_{m},\dot{x}_{m},g_{m}\}$) While satisfying the physical equation of motion in equation 2, e.g. $f\rightarrow 0$. Here $W_{\theta}$$b_{\theta}$ are weights and biases of the neural network. $J_{D}(\theta)$ is the data loss based on the measurements and $J_{P}(\theta)$ denotes the physics loss, which imposes a physical constraint on the neural network.
2.1 1D U-net
Figure 2: Proposed 1D version of U-net: The U-Net architecture is mainly divided into four major components: the encoder blocks, bottleneck, decoder blocks, and the output convolution block. The encoder blocks perform convolution and pooling operations, generating skip connections and providing input to subsequent encoder blocks after pooling. The decoder blocks involve upsampling, followed by concatenation with the skip connections from the corresponding encoder block outputs, and subsequent convolution operations. The bottleneck layer, serving as a bridge, consists solely of a convolution block, linking the encoder and decoder paths. The output from the final decoder block is fed to the output convolution block, where it first passes through a Conv1D layer with a number of filters equal to the number of channels required in the final output (kernel size = 1), and sigmoid activation, followed by a second Conv1D layer and a linear activation function, which further processes the output to format the final output as a three-dimensional array, where the entries are sampled in the first dimension, time history steps in the second dimension, and output features in the last dimension. Here, ‘ConvBlock’is a fundamental building block consisting of a convolutional layer, followed by Batch Normalisation, and a ReLU activation function.
U-Net [ronneberger2015u] is an established framework originally developed for biomedical image segmentation [unet1, unet2]. It has a symmetrical U-shaped structure, consisting of an encoder and a decoder connected by skip connections. The encoder captures contextual information by progressively downsampling the input image, while the decoder reconstructs the image by upsampling and combining high-resolution features from the encoder, ensuring precise localisation.
The ‘Slow Feature Analysis’ hypothesis [wiskott2002slow] posits that meaningful features in dynamic signals often evolve slowly over time. Considering the encoder-decoder components and the ability to capture contextual information and combine resolution features, U-nets are already doing it, i.e. computing features at different time scales with two-dimensional convolutions and combining them to make predictions at the same resolution as the input. One such adaptation is the one-dimensional U-Net, suitable for time-series analysis [stoller2019sequnetonedimensionalcausalunet]. In this study, we propose a causal 1D U-Net to process seismic time-series data that captures such features across multiple temporal scales while maintaining causal structure, ensuring outputs at a given time step depend only on current and past inputs. This enhances the model’s suitability for sequential tasks and integration with LSTM networks.
The architecture comprises four parts: the encoder (contractive path), the bottleneck, the decoder part (expansive path) and finally, the output convolution block. The input is processed sequentially through two encoder blocks, having 50 and 100 filters respectively. Each encoder block generates two outputs: one is a skip connection (or shortcut connection) for the respective decoder block, and the other (output after 1D pooling operation) serves as the input for the subsequent encoder block. The final encoder block provides inputs for the bottleneck layer and also creates a skip connection. The encoder block comprises two sequential ‘ConvBlock’ units. Each ‘ConvBlock’ executes a series of operations: a 1D convolution ( kernel size = 2, padding = ‘causal’ ), followed by Batch Normalization, and finally a ReLU activation function.
The resulting output from this process serves as both a skip connection (shortcut path) and as the input for the subsequent MaxPooling1D operation. MaxPooling1D is a down-sampling operation typically used in convolutional neural networks (CNNs). This method reduces the dimensionality of the input tensor by extracting the maximum value from within a sliding window, known as the pool size (set to 2), across the input. So, the encoder block consists of two main components: the convolution blocks and the max pooling block.
For down-sampling, the ‘Conv1D’ operation with strides=2 (padding= ‘causal’) can also be implemented, but here in this case ‘MaxPooling1D’ is used to emphasize keeping feature extraction and down-sampling as distinct operations. And to ensure that the most significant features (e.g. sudden large significant peaks in ground acceleration data) are preserved during downsampling. There are also a few additional advantages. It is more straightforward than implementing ‘Conv1D’ with a stride of 2, and it reduces feature map size without adding extra learnable parameters.
The bottleneck layer also comprises two sequential ’ConvBlock’ units, using 200 filters without any pooling operation. This layer acts as a crucial bridge between the encoder and decoder, capturing complex, high-level features while maintaining the spatial dimensions of the feature maps. In the decoder stage, the first decoder block receives the output from the bottleneck layer, along with the skip connection from the second encoder block, and uses these inputs to upsample the feature map.
Within the decoder block, a 1D upsampling operation is performed on the input tensor that repeats each temporal step size times along the time axis. In this case, as the parameter size=2 is specified, the layer performs upsampling by repeating each temporal step of the input sequence twice along the time axis. This step helps to gradually reconstruct the original input size. After the 1D upsampling operation, the feature map is concatenated with the corresponding skip connection from the encoder blocks. Then the output passes through two ConvBlocks. This sequence is repeated in each of the two decoder blocks, with 100 and 50 filters, respectively, to progressively refine the feature maps and reconstruct the final output with high accuracy and detail. The output from the final decoder block is then first passed through a Conv1D layer with a number of filters equal to the number of channels required in the final output, a kernel size of 1, and sigmoid activation, which reduces the channels to match the dimension of the final output required and maps the output values between 0 and 1. The output has the same length as the input due to padding. This is followed by a second Conv1D layer and a linear activation function, which further processes the output by expanding the channels to format the final output as a three-dimensional array, where the entries are sampled in the first dimension, time history steps in the second dimension, and output features in the last dimension, this is very crucial otherwise the deep LSTM network cannot be implemented subsequently. The network architecture is shown in Figure 2 and represented in the table 1. The U-Net output feeds directly into the LSTM module.
Layer Type Number of Filters Kernel Size Stride Activation
Encoder 1 Conv Block (x2) 50 2 1 ReLU
Encoder 1 Max Pooling - 2 2 -
Encoder 2 Conv Block (x2) 100 2 1 ReLU
Encoder 2 Max Pooling - 2 2 -
Bottleneck Conv Block (x2) 200 2 1 ReLU
Decoder 1 UpSampling1D - - - -
Decoder 1 Conv Block (x2) 100 2 1 ReLU
Decoder 2 UpSampling1D - - - -
Decoder 2 Conv Block (x2) 50 2 1 ReLU
Output Convolutional Number of output channels 1 1 Sigmoid
Final output Convolutional Number of output channels 1 1 Linear
Table 1: Proposed 1D U-Net Architecture. Note: Inside each Conv Block (Convolution Block), a consecutive set of the following operations: Convolution → Batch Normalization → ReLU activation takes place.
2.2 Deep LSTM Network
Each LSTM layer consists of a set of LSTM cells, as depicted in Figure 3. Like the neural nodes in traditional ANNs, each LSTM cell contains an independent set of weights and biases shared across the entire temporal sequence within the layer. First, the basic architecture of the deep LSTM network for sequence-to-sequence modeling is presented[zhang2019deep]. The architecture includes a number of hidden layers - additional to input and output layers - as shown in Figure 3(a), some of which are LSTM layers, while others are fully connected layers. The LSTM cell includes four components: an internal cell, an input gate, a forget gate, and an output gate. The internal cell remembers the cell state from the previous time step via a self-recurrent connection. The input gate maintains the flow of input activations into the internal cell state, while the output gate controls the flow of output activations from the LSTM cell. The forget gate scales the internal cell state, allowing the LSTM cell to adaptively forget or reset its memory. Through the coordinated actions of the input, forget, and output gates, the cell state can selectively transmit essential information along the temporal sequence, capturing both long-term and short-term dependencies in a dynamic system effectively.
Figure 3: Schematic of deep LSTM networks: (a) corresponds to the network architecture of a deep LSTM network featuring $m$ LSTM layers and multiple fully-connected layers for sequence-to-sequence modeling and (b) corresponds to the architecture of a typical LSTM cell at the $l$th layer and time $t$, showing cell input $X^{(l)}_{t}$, cell output $Y^{(l)}_{t}$, cell state $c^{(l)}_{t}$, hidden state $h^{(l)}_{t}$, and gate variables $\{f^{(l)}_{t},i^{(l)}_{t},\tilde{c}^{(l)}_{t},o^{(l)}_{t}\}$, respectively.
Let us consider at time step t (where ( $t=1,2,3,\ldots,n$ ); with n being the total number of time steps) and within the $l$-th LSTM network layer, the input state to the LSTM cell as $x_{t}^{(l)}$, and the forget gate as $f_{t}^{(l)}$, the input gate as $i_{t}^{(l)}$, the output gate as $o_{t}^{(l)}$, the cell state memory as $c_{t}^{(l)}$ and hidden state memory as $h_{t}^{(l)}$. Similarly, for the previous time step t-1, we denote the cell state memory as $c_{(t-1)}^{(l)}$ and hidden state memory as $h_{(t-1)}^{(l)}$. The relationship between these variables can be expressed as follows:
$f_{t}^{(l)}=\sigma\left(W_{xf}^{(l)}\cdot x_{t}+W_{hf}^{(l)}\cdot h_{t-1}^{(l)}+b_{f}^{(l)}\right)$ (3)
$i_{t}^{(l)}=\sigma\left(W_{xi}^{(l)}\cdot x_{t}+W_{hi}^{(l)}\cdot h_{t-1}^{(l)}+b_{i}^{(l)}\right)$ (4)
$\tilde{c}_{t}^{(l)}=\tanh\left(W_{xc}^{(l)}\cdot x_{t}+W_{hc}^{(l)}\cdot h_{t-1}^{(l)}+b_{c}^{(l)}\right)$ (5)
$o_{t}^{(l)}=\sigma\left(W_{xo}^{(l)}\cdot x_{t}+W_{ho}^{(l)}\cdot h_{t-1}^{(l)}+b_{o}^{(l)}\right)$ (6)
$c_{t}^{(l)}=f_{t}^{(l)}\odot c_{t-1}^{(l)}+i_{t}^{(l)}\odot\tilde{c}_{t}^{(l)}$ (7)
$h_{t}^{(l)}=o_{t}^{(l)}\odot\tanh\left(c_{t}^{(l)}\right)$ (8)
Where $W_{(\alpha\beta)}^{(l)}$ (with $\alpha$ = $\{x,h\}$} and $\beta={\{f,i,c,o\}}$ ) denotes weight matrix corresponding to different inputs (e.g. $x_{t}^{l}$ or $h_{t}^{l}$ ) different gates (e.g., input gate, forget gate, tanh layer, or output gate). And $b_{\beta}^{(l)}$ represents the corresponding bias vectors; the lth layer of the LSTM network is denoted by the superscript l. Here, $\tilde{c}_{t}^{(l)}$ denotes a vector of intermediate candidate values created by a tanh layer shown in Fig. 1(b); $\sigma$ is the logistic sigmoid function; tanh is the hyperbolic tangent function; $\odot$ denotes the Hadamard product (elementwise product). The complex connection system within each LSTM cell makes the deep LSTM network powerful in sequence modelling, in which the fully connected layers are beneficial to map the temporal feature maps to the corresponding output space.
The output of the final dense layer produces a time-series prediction of the state-space variables $z(t)={x(t),\dot{x}(t),g(t)}$, where $x$ is displacement, $\dot{x}$ is velocity, and $g$ is restoring force. While this formulation accurately captures temporal patterns, it lacks explicit knowledge of the governing physical laws.To address this limitation, the predicted outputs are further processed using a graph-based tensor differentiator to compute their temporal derivatives. A specialized loss function is then formulated, tailored to the specific requirements of each scenario. The tensor differentiator and the loss function are described in detail in Sections 2.3 and  2.4, respectively. Training proceeds by minimizing the loss:
$\hat{\theta}\coloneqq\arg\min_{\theta}J(\theta)$ (9)
The model is implemented in Python using Keras [chollet2018keras] with TensorFlow backend. Training is conducted on Google Colab Pro with GPU acceleration. All hyperparameters, including the number of LSTM layers $m$, number of hidden units per layer, batch size, learning rate, and optimizer settings, are specified in the Supplementary Information.The detailed network architecture of the proposed framework is represented in Figure 1.
2.3 Graph Based Tensor Differentiator
Numerical differentiation is an essential method widely used in computational mathematics, particularly in solving differential equations and gradient-based optimization. Finite differences (FD) approximate derivatives by combining nearby function values using a set of weights. Here in this study to calculate the derivatives of the state space outputs of the neural network a graph-based tensor differentiator utilizing finite difference approach is implemented.[phycnn] This section gives detailed explanation about the construction of finite difference matrix ($\Phi_{t}]$) to calculate the first order derivatives providing second order accuracy meaning the truncation error is of order 2 (i.e. $O(\Delta t^{2})$). For all the interior points central difference stencil is utilized whereas for the first and last point forward and backward difference stencil is implemented.
Let us assume a function $u=[u_{0},u_{1},\ldots,u_{n-1}]$ is sampled at time $t_{0},t_{1},\ldots,t_{n-1}$. Here the data $u_{i}=u(t_{i})$ given at equispaced points $t_{i}=t_{0}+i\Delta t$, for $i=0,1,2,\ldots,n-1$ where $\Delta t$ is the step size. The first order derivatives are approximated as:
$\left.\frac{du}{dt}\right|_{t_{i}}\approx\begin{cases}\displaystyle\frac{-\frac{3}{2}u_{i}+2u_{i+1}-\frac{1}{2}u_{i+2}}{\Delta t},&i=0\\ \displaystyle\frac{u_{i+1}-u_{i-1}}{2\Delta t},&0<i<n-1\\ \displaystyle\frac{\frac{1}{2}u_{i-2}-2u_{i-1}+\frac{3}{2}u_{i}}{\Delta t},&i=n-1\end{cases}$ (10)
In order to compute the first-order derivative of vector $\{u\}=[u_{0},u_{1},\ldots,u_{n-1}]^{T}$ as mentioned in equation 10, which is $\{u_{t}\}$ , matrix $\Phi_{t}]$ is constructed which transforms $\{u\}$ to $\{u_{t}\}$.
$\{u_{t}\}=[\Phi_{t}]\{u\}$ (11)
Where $\Phi_{t}]$ is an $n\times n$ finite difference matrix whose entries constitute the optimal weights for approximating first derivatives at each discrete point through a weighted combination of neighboring function values. The complete finite difference matrix can be expressed as:
$\Phi_{t}=\frac{1}{\Delta t}\begin{bmatrix}-3/2&2&-1/2&0&\ldots&0\\ -1/2&0&1/2&0&\ldots&0\\ 0&-1/2&0&1/2&\ldots&0\\ \vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ 0&\ldots&0&1/2&-2&3/2\end{bmatrix}_{n\times n}$
2.4 Loss function
In this study, three different loss functions are utilised for each of the two numerical validation scenario and the experimental validation of the model. To incorporate physical constraints, temporal derivatives $\dot{z}(t)={x_{t}(t),\dot{x}_{t}(t),g_{t}(t)}$ are estimated using a graph-based differentiator [zhang2019deep], which implements finite-difference approximations over a local stencil. This component enables embedding of physics-based residuals into the loss function. For the first scenario of the numerical validation where full-state measurements are available for training, the total loss function used for training combines data loss and physics-based loss:
$J(\theta)=w_{1}\cdot J_{D}(\theta)+w_{2}\cdot J_{P}(\theta)$ (12)
$J_{D}(\theta)=\frac{1}{N}\left\|x^{p}-x^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}^{p}-\dot{x}^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|g^{p}-g^{m}\right\|_{2}^{2}$ (13)
$J_{P}(\theta)=\frac{1}{N}\left\|\dot{x}^{p}-x_{t}^{p}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}+g^{p}+\Gamma\ddot{x}_{g}\right\|_{2}^{2}$ (14)
Where $J_{D}(\theta)$ is the data loss based on the measurements and $J_{P}(\theta)$ denotes the physics loss . Superscripts $p$ and $m$ denote predicted and measured quantities, respectively. When certain measurement channels are unavailable, corresponding loss terms are omitted. In the physics loss the second term incorporates the physical laws in terms of the governing equation where $\dot{x}_{t}^{p}$ represents the derivative of the predicted velocity (which corresponds to the acceleration). The first term serves as a consistency or sanity check, ensuring that the predicted state-space variables align with their physical definitions. Specifically, the predicted velocity $\dot{x}^{p}$ should be equal with the derivative of the predicted displacement($x_{t}^{p}$) computed using the graph-based tensor differentiator. In this study, equal weights ($w_{1}=w_{2}=1$) are used, but these can be tuned depending on the application. For the second case of numerical validation, although the system information—such as the mass matrix, stiffness, and damping coefficients—is available, only acceleration measurements are used as training data. A modified loss function is architected to incorporate the governing equations and enable training using acceleration data alone to accommodate this limitation. The formulation of this loss function is presented below:
$J(\theta)=\frac{1}{N}\left\|\dot{x}^{p}-x_{t}^{p}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}-\ddot{x}^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}+g^{p}+\Gamma\ddot{x}_{g}\right\|_{2}^{2},$ (15)
This modified loss function’s first term is identical to the original physics loss’s first term, ensuring consistency between predicted state variables. The third term also remains unchanged from the second term of the original physics loss, enforcing adherence to the physical laws. In the second term the $\dot{x}_{t}^{p}$ represents the time derivative of the predicted velocity, computed via the graph-based tensor differentiator, and enforced to match the available measured acceleration data.
For the experimental validation where we dont have a complete description of the structural system—including mass, stiffness, damping properties, and interaction effect, it is not possible to implement physics based loss term and the model relies solely on the data loss. In this purely data-driven scenario, only acceleration data gathered from the sensors are available for training. The model is designed to predict displacements, which are then differentiated twice using the graph-based tensor differentiator to obtain the corresponding accelerations ($x_{tt}$) using tensor differentiator. The loss function implemented for this case is presented below:
$J(\theta)=\frac{1}{N}\left\|\ddot{x}^{m}-x_{tt}^{p}\right\|_{2}^{2}$ (16)
2.5 Numerical Validation:
For numerical validation, we consider a nonlinear single-degree-of-freedom system subjected to ground motion excitation. The system is governed by
$m\ddot{x}+\underbrace{c\dot{x}+k_{1}x+k_{2}x^{3}}_{\text{h}}=-m\Gamma\ddot{x}_{g}$ (17)
where $m=1\,\mathrm{kg}$, $c=1\,\mathrm{Ns/m}$, $k_{1}=20\,\mathrm{N/m}$, and $k_{2}=200\,\mathrm{N/m}$. The restoring force $h$ is normalized by mass to yield $g=h/m$, following Zhang [phycnn]. Two scenarios are considered. In the first, full-state measurements—displacement, velocity, and restoring force—are assumed available for training. These can be extracted from numerical simulations, with $g$ computed via $g=-\ddot{x}-\Gamma\ddot{x}_{g}$. In the second, only acceleration data are used, representing a more constrained condition. This setup enables direct assessment of the proposed model’s robustness under varying levels of data availability.
2.5.1 Case 1: Full-state measurements available for training
We consider the scenario in which full-state measurements—displacement ($x$), velocity ($\dot{x}$), and restoring force (gg)—are available for training the model. Following Zhang [phycnn], we adopt a dataset of 100 numerical simulations of a nonlinear single-degree-of-freedom system subjected to synthetic earthquake ground motions drawn from the PEER Strong Motion Database [chiou2008nga], representing a 10% probability of exceedance in 50 years. Each simulation spans 50 seconds with a sampling rate of 20 Hz, resulting in 1001 time steps per example. Out of these, 10 records are randomly selected for training, while the remaining 90 are reserved for validation. The proposed PhyULSTM framework (Fig. 1) is trained to map ground motion acceleration to system response using these data. The model input and output tensors are of shapes [10, 1001, 1] and [10, 1001, 3], respectively. Training follows the architecture and loss function detailed in 2.4.
$J(\theta)=w_{1}\cdot J_{D}(\theta)+w_{2}\cdot J_{P}(\theta)$
$J_{D}(\theta)=\frac{1}{N}\left\|x^{p}-x^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}^{p}-\dot{x}^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|g^{p}-g^{m}\right\|_{2}^{2}$
$J_{P}(\theta)=\frac{1}{N}\left\|\dot{x}^{p}-x_{t}^{p}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}+g^{p}+\Gamma\ddot{x}_{g}\right\|_{2}^{2}$
Figure 4 illustrates the predicted displacement response for three unseen ground motion using both PhyULSTM and PhyCNN models. The proposed PhyULSTM architecture demonstrates excellent predictive performance, accurately reproducing the displacement time history with close agreement to the ground truth. Notably, PhyULSTM captures both the overall trend and the rapid fluctuations in the response that are significantly underrepresented by PhyCNN. This difference is especially prominent in regions with sharp nonlinear behavior, underscoring PhyULSTM’s superior ability to generalize across previously unseen inputs. Figure 5 further compares the predicted velocity and nonlinear restoring force for the a representative ground motion. Again, PhyULSTM closely matches the true values obtained from direct numerical integration of the governing equations. In contrast, PhyCNN tends to smooth out higher-frequency variations and fails to capture abrupt transitions in the restoring force dynamics. This limitation becomes particularly evident in the high-gradient regions, where PhyULSTM maintains fidelity to the true signal.
The nonlinear dynamic behavior of the system is further examined through hysteresis plots of the restoring force versus displacement and velocity, as shown in Fig. 6. These plots serve as critical indicators of the model’s ability to capture complex energy dissipation mechanisms and stiffness degradation inherent in nonlinear structural systems.
Figure 4: Time histories of displacement responses predicted by PhyULSTM (red, left column) and PhyCNN (red, right column) compared against the reference solution (blue). Each row corresponds to a different earthquake input (1, 2, and 3). Subplots are labeled 1a, 1b, 2a, 2b, 3a, and 3b for reference, where the left subplots (a) show PhyULSTM predictions and the right subplots (b) show PhyCNN predictions. Across all three test cases, PhyULSTM consistently tracks the true system behavior with higher accuracy, capturing both transient peaks and steady-state oscillations. In contrast, PhyCNN fails to reproduce finer-scale dynamics and underestimates displacement amplitudes during critical phases.
The results in Fig. 6 thus confirm that PhyULSTM is not only capable of accurate time-series predictions but also preserves the nonlinear physics of the underlying system, making it a reliable tool for surrogate modeling in structural dynamics.
Figure 5: Predicted time histories of velocity ($\dot{x}$) and nonlinear restoring force ($g$) for a representative ground motion are shown.The left column shows the responses predicted by PhyULSTM, and the right column shows those predicted by PhyCNN. Subplots are labeled 4a, 4b, 4c, and 4d for reference: 4a and 4c correspond to velocity and normalized restoring force predicted by PhyULSTM, while 4b and 4d correspond to the same predicted by PhyCNN. The PhyULSTM model captures both quantities with high fidelity compared to the numerical reference, successfully reproducing subtle fluctuations and transient peaks.
Figure 6: PhyULSTM accurately replicates the mechanical response of the nonlinear system, as evidenced by the hysteresis plots of restoring force $g$ versus displacement $x$ (left) and velocity $\dot{x}$ (right). The model successfully captures the characteristic loops, amplitude variations, and shape distortions typical of nonlinear restoring forces, affirming its ability to infer high-order dynamic behavior from limited observations.
Figure 7 presents the results of a regression analysis comparing the predicted and true displacement responses for both PhyULSTM and PhyCNN models. The proposed PhyULSTM framework demonstrates remarkable predictive fidelity, achieving a maximum correlation coefficient of 0.998 and a minimum of 0.83 across all 90 unseen test cases. These values indicate near-perfect agreement in most scenarios, even under conditions involving significant nonlinearity and variability in input ground motions. In contrast, the PhyCNN model exhibits noticeably lower predictive accuracy, with correlation coefficients ranging from 0.97 down to 0.577, suggesting poorer generalization and a tendency to underperform in more complex test cases.
Figure 7: Histogram of predicted versus ground-truth displacements for PhyULSTM (left) and PhyCNN (right). PhyULSTM achieves a maximum correlation coefficient of 0.998 and a minimum of 0.83 across all test cases, indicating robust accuracy and generalization. In contrast, PhyCNN exhibits lower performance, with correlation values ranging from 0.97 to 0.577. These results highlight the superior learning capacity of the proposed physics-informed LSTM architecture.
Additionally, in the case of the PhyCNN model, 61.1% of the predicted displacement responses exhibit a Pearson correlation coefficient greater than 0.9 with the reference values. In contrast, the proposed PhyULSTM model achieves a significantly improved performance, with 96.7% of its predicted displacements exceeding the 0.9 correlation threshold. This comparison underscores the robustness and reliability of PhyULSTM, particularly in scenarios with limited training data. The ability of PhyULSTM to maintain high prediction accuracy across diverse and challenging test inputs highlights its superior generalization capacity and establishes it as a more effective surrogate modeling framework for nonlinear dynamic systems.
2.5.2 Case 2: Training with acceleration measurements only
Figure 8: Time histories of displacement responses predicted (red) by PhyULSTM (left) and PhyCNN (right) are compared against reference numerical solutions (blue) obtained via direct integration. Each row corresponds to a different earthquake input (1, 2, and 3). Subplots are labeled 1a, 1b, 2a, 2b, 3a, and 3b for reference, where the left subplots (a) show PhyULSTM predictions and the right subplots (b) show PhyCNN predictions. PhyULSTM consistently captures transient dynamics and amplitude variations with high fidelity across three distinct seismic events, whereas PhyCNN exhibits noticeable deviations, especially in peak displacement magnitudes and oscillation patterns.
In practical structural monitoring, accelerometers are often the only sensors installed, providing solely acceleration time histories. Traditional deep learning models approaches attempt to infer displacements by numerical integration of accelerations before training predictive models. However, this introduces significant numerical errors that degrade prediction accuracy. In contrast, the proposed PhyULSTM framework directly incorporates physics constraints into its training, enabling accurate displacement field predictions using only acceleration inputs. Leveraging a specialized 1D U-net architecture combined with LSTM units, the model extracts salient features from acceleration signals to infer the full state vector ${z}(t)=\{x(t),\dot{x}(t),g(t)\}$. These predictions are further refined by a tensor differentiator employing central finite differences to compute derivatives $\dot{z}(t)=\{x_{t}(t),\dot{x}_{t}(t),g_{t}(t)\}$. The loss function is designed to penalize deviations between predicted derivatives and measured accelerations, enforcing physical consistency as
$J(\theta)=\frac{1}{N}\left\|\dot{x}^{p}-x_{t}^{p}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}-\ddot{x}^{m}\right\|_{2}^{2}+\frac{1}{N}\left\|\dot{x}_{t}^{p}+g^{p}+\Gamma\ddot{x}_{g}\right\|_{2}^{2},$
where only acceleration measurements $\ddot{x}^{m}$ contribute to the loss due to data availability. The model is trained on 50 randomly selected datasets and tested on 50 unseen datasets.
Figure 8 illustrates the displacement time histories predicted by the proposed PhyULSTM model alongside those from the baseline PhyCNN and the ground truth numerical solutions for three distinct, previously unseen seismic excitation inputs. The PhyULSTM predictions (left) closely track the true system response (blue), accurately capturing both the amplitude and phase of complex oscillatory behavior inherent to nonlinear structural dynamics. In contrast, the PhyCNN predictions (right) exhibit noticeable deviations, particularly during peak displacement intervals and rapid transient phases, indicating an underestimation of key dynamic features. Figure 9 further highlights the predictive fidelity of PhyULSTM by comparing the velocity ($\dot{x}$) and nonlinear restoring force ($g$) time histories for a representative unseen seismic event. The model faithfully reproduces the intricate temporal variations and sharp fluctuations characteristic of nonlinear hysteretic behavior, which are only partially captured by PhyCNN.
Figure 9: PhyULSTM accurately reproduces velocity ($\dot{x}$) and restoring force ($g$) time histories, closely matching the ground truth obtained from numerical integration. The left column shows the responses predicted by PhyULSTM, and the right column shows those predicted by PhyCNN. Subplots are labeled 4a, 4b, 4c, and 4d for reference: 4a and 4c correspond to velocity and normalized restoring force predicted by PhyULSTM, while 4b and 4d correspond to the same predicted by PhyCNN. The model effectively captures the nonlinear system response including sharp fluctuations, which alternative methods fail to capture effectively.
Figure 10: Correlation between predicted and true displacements across all test cases is higher for PhyULSTM (left), with maximum and minimum correlation coefficients of 0.996 and 0.764 respectively, compared to PhyCNN (right) which attains 0.984 and 0.727. This demonstrates the superior generalization and predictive robustness of PhyULSTM when trained solely on acceleration data.
The regression analysis summarised in Figure 10 quantitatively confirms these observations. PhyULSTM achieves a maximum correlation coefficient of 0.996 and maintains a minimum of 0.764 across all test cases, demonstrating robust generalization even when trained exclusively with acceleration measurements. By contrast, PhyCNN’s correlation coefficients span a lower range (0.984 maximum to 0.727 minimum), underscoring its comparatively weaker predictive accuracy and generalization. Additionally, in the case of the PhyCNN model, 80% of the predicted displacement responses exhibit a Pearson correlation coefficient greater than 0.9 with the reference values. In contrast, the proposed PhyULSTM model achieves a significantly improved performance, with 94% of its predicted displacements exceeding the 0.9 correlation threshold. These comprehensive results establish that PhyULSTM is a powerful tool for real-world structural health monitoring and seismic response analysis.
2.6 Experimental Validation
Structural systems are inherently complex, often characterized by nonlinear behavior, heterogeneous materials, and uncertain boundary conditions. In many real-world scenarios, obtaining a complete description of the structural system—including mass, stiffness, damping properties, and interaction effects—is infeasible due to limited instrumentation, inaccessible subsystems, or cost constraints. Traditional numerical methods, while powerful, rely on prior knowledge of the governing equations and boundary conditions. As such, they are generally limited to idealized models and may exhibit significant deviations from actual system behavior when applied to partially known or ill-defined systems. To overcome these limitations, we extend the PhyULSTM framework—originally formulated with embedded physics-based constraints—to operate under data-driven settings where physical knowledge of the system is incomplete or entirely unavailable. In this formulation, the loss function excludes physics-based residual terms, relying solely on measurement-based error minimization. The network architecture remains unchanged, retaining the hybrid encoder–decoder configuration that combines a one-dimensional U-Net with a Long Short-Term Memory (LSTM) decoder . This architecture is capable of learning complex mappings from input ground motions to structural responses, even in the absence of explicit physical regularization.
To validate this data-driven variant of the PhyULSTM model, we utilize recorded seismic response data from a six-story reinforced concrete hotel building located in San Bernardino, California. The structure, designed in 1970, has been instrumented with nine accelerometers on the first, third, and roof levels in both principal directions. The dataset, curated by the Center for Engineering Strong Motion Data (CESMD) [haddadi2012report], comprises earthquake recordings collected between 1987 and 2018. These events span a broad range of magnitudes and frequency content, thereby providing a representative and challenging benchmark for data-driven structural response modeling. For consistency and reproducibility, the dataset partitioning follows the protocol established in Zhang et al. [phycnn], which involves stratified sampling into training, validation, and test subsets. The training objective for the data-driven PhyULSTM model is to learn a mapping from ground acceleration inputs to structural displacement outputs.
Figure 11: Predicted displacement response at the 3rd floor (DOF-1) of the six-story instrumented hotel building in San Bernardino, California, subjected to the 2014 Big Bear Lake and 2016 Loma Linda earthquakes. The model was trained using historical seismic data from 1987–2018, relying solely on acceleration measurements from strategically placed sensors, without incorporating any explicit physical constraints in the loss function. Despite this, the PhyULSTM model successfully captures the dynamic behavior of the structure, including transient displacement patterns and resonant oscillations, aligning well with sensor-derived ground truth.
The model is trained by minimizing the discrepancy between predicted and measured accelerations, defined by the following objective function:
$J(\theta)=\frac{1}{N}\left\|\ddot{x}^{m}-x_{tt}^{p}\right\|_{2}^{2}$
where $\ddot{x}^{m}$ denotes the measured structural acceleration and $x_{tt}^{p}$ represents the predicted acceleration obtained by applying a central finite difference operator to the predicted displacement time series. The network is trained using 11 earthquake records, with 4 events used for validation. Once trained, the model is tested on previously unseen events to evaluate its generalization capability. Figure 11 and Figure 12 present the predicted displacement responses at the 3rd floor and roof levels, respectively, of the six-story instrumented hotel building in San Bernardino, California, subjected to the 2014 Big Bear Lake and 2016 Loma Linda earthquakes. The PhyULSTM model was trained using only acceleration time histories obtained from sensors distributed across the lower, middle, and upper stories of the structure, without incorporating explicit physics-based constraints in the loss function. Despite this, the model accurately captures essential structural behaviors across events of varying intensity, duration, and frequency content. These findings highlight the robustness and generalization capacity of the PhyULSTM architecture for modeling real-world nonlinear structural systems under seismic excitation. Importantly, the model achieves this without requiring detailed system identification or constitutive modeling, positioning it as a viable surrogate modeling tool for structural health monitoring applications, particularly in scenarios where only partial or noisy sensor data are available.
Figure 12: Predicted roof-level displacement (DOF-2) of the six-story San Bernardino hotel building under the 2014 Big Bear Lake and 2016 Loma Linda earthquake events.
3 Conclusion
This study introduces a hybrid deep learning framework, PhyULSTM, that integrates the architectural strengths of U-net and LSTM within a physics-informed training regime for the surrogate modeling of nonlinear structural systems under seismic excitation. When available, governing physical laws are incorporated directly into the loss function to guide learning, mitigate overfitting, and enhance generalization, particularly in data-scarce environments. Moreover, the model outperforms existing seismic response prediction models like PhyCNN and is also capable of operating without explicit physical constraints, enabling its application to systems with incomplete or unknown constitutive characteristics. Extensive numerical experiments on nonlinear single-degree-of-freedom systems and experimental validation using sensor data from a mid-rise concrete building in San Bernardino demonstrate that the PhyULSTM framework outperforms existing physics-informed CNN architectures. It exhibits superior predictive accuracy, higher correlation in regression analysis, and robust performance under varied excitation characteristics, including amplitude and frequency content. These results underscore the framework’s ability to generalize structural response behavior under seismic loading, even when trained on limited and noisy data.
The modular nature of the PhyULSTM model allows it to be extended to other structural typologies, including bridges, tunnels, and other civil infrastructure. Furthermore, the combination of U-net and LSTM makes the architecture well-suited for broader time-series modeling tasks beyond structural engineering, such as applications in signal processing and sequential decision-making. Overall, this study establishes PhyULSTM as a reliable, interpretable, and physically grounded metamodeling tool that advances the state of surrogate modeling for structural dynamics and provides a foundation for future developments in hybrid physics-informed machine learning.